Steve wrote a x 1 = a. He said that the equation is true of any value of a. Do you agree with Steve? Tell why or why not.

Is it a reasonable solution or not? 85 / 5= r; r =7
1
Expert's answer
2012-09-25T11:04:43-0400
Steve is right. Any real value of a satisfy this equation. Let's derive this equation by a:

& (a x 1) / a = a / a
&
& 1 = 1
We've got identity.

&
Is it a reasonable solution or not? 85 / 5= r; r = 7.

No, the reasonable solution is r = 17.